Combustion of hydrocarbons such as methane (CH4) produces carbon dioxide, a greenhouse gas. Greenhouse gases in the Earths
Answer #1

1)

The reaction is:

CH4(g) + 2 O2(g) —> CO2(g) + 2 H2O(g)

2)

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass of CH4 = 4.2*10^2 g

mol of CH4 = (mass)/(molar mass)

= 4.2*10^2/16.04

= 26.18 mol

According to balanced equation

mol of CO2 formed = moles of CH4

= 26.18 mol

Given:

P = 1.0 atm

n = 26.18 mol

T = 17.0 oC

= (17.0+273) K

= 290 K

use:

P * V = n*R*T

1 atm * V = 26.18 mol* 0.08206 atm.L/mol.K * 290 K

V = 623 L

Answer: 623 L